OpenStudy (anonymous):

fundamental question. 17.2=nX25.5 to find n i could either divide 25.5 on both sides to get 17.2/25.5 but would i be wrong if i divided n by sides to get 17.2/n=25.5 then times 1/17.2 on bothsides to get 25.5/17.2 HOW IS THIS WRONG BUT The OTHER WAY IS RIGht?!

OpenStudy (anonymous):

17.2=n*25.5 sorry that last equation might have been confusing with the X being times.

OpenStudy (anonymous):

In both the cases u get the same value

OpenStudy (anonymous):

no u dont...

OpenStudy (anonymous):

17.2/25.5= .67

OpenStudy (anonymous):

25.5/17.2 = 1.4 something something

OpenStudy (anonymous):

1/n = 25.5/17.2 in the later case as per ur question

OpenStudy (anonymous):

OOOH damn u smart....

OpenStudy (anonymous):

i feel dumb now.. thanks for the education.

OpenStudy (anonymous):

to get n it should be 1/1.4 which is 0.714